OpenStudy (anonymous):

put this in point slope form.. (y-y1)=m(x-x1) .. x1=-1 f(x1)=3 x2= -4 f(x2) = 2 @rebeccaskell94

OpenStudy (anonymous):

I don't see why you do this yourself, you have two distinct points, just use the slope formula

OpenStudy (anonymous):

i dont know how to do f(x)! i dont know what that is

OpenStudy (anonymous):

F(x) is just your y

OpenStudy (anonymous):

ohhh thats dumb why cant they just say y?

OpenStudy (lgbasallote):

im still waiting for @rebeccaskell94 to answe this :S

OpenStudy (anonymous):

She cant

OpenStudy (anonymous):

It's to "confusing".......lol

OpenStudy (lgbasallote):

main point is you cant use point slope...you use two point formula

OpenStudy (anonymous):

( y+3)=1/5(x+1) is that right????

OpenStudy (anonymous):

@NightShade am i right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you have two points, \((-4,2)\) and \((-1,3)\)

OpenStudy (anonymous):

slope is \[\frac{3-2}{-1-(-4)}=\frac{1}{-1+4}=\frac{1}{3}\]